Two long parallel wire are 10cm carrying a current of 6A and 4A. Calculate the force of acting on a 100cm length of

1
Expert's answer
2022-01-27T13:25:07-0500

"F=\frac{μ_02I_1I_2l}{4\pi d}"

Where

"I_1=6A"

"I_2=4A"

"d=10\u00d710^{-2}M =0.1M"

"l=100\u00d710^{-2}M =1M"

"μ_0=4\pi×10{^-7}"


"F=\frac{4\pi×10^{-7}×2×6×4×1}{4\pi ×0.1}=4.8×10^{-5}N"